what transition of Li2+ spectrum will have same wavelength as that of second line of balmer series in he+ spectrum.

Grade:11

2 Answers

Umakant biswal
7 years ago
@ rishant 
can you recall the following formula ?? 
lambda=1/z^2{(1/n1^2)-(1/n2^2)},  
if u can then , u can surely , solve this question . "
BUT THIS QUESTION IS WRONG , SOME DATA GIVEN HERE IS WRONG . 
LET ME TO EXPLAIN WHY .. 
AFTER applying the above formula 
so we have 1/9(1/n1^2-1/n2^2)=1/4(1/4-1/6) 
so, after that we have by solving 
 1/n1^2-1/n2^2=9/16-9/64
and bysolving further i am getting 27/64 . 
we have to write 27/64 in the way of 1/n1^2-1/n^2
and thats not possible anyhow . 
SO, THE QUESTION IS WRONG 
KINDLY RECHECK THE QUESTION PLZZ.. 
ALL THE BEST ..
 
monu kumar
7 years ago
their is some mistakes in the question even you try u can use the formula of rybergs and try  to solve it but it cannot be solved because the question has an error